# Selmer stability for elliptic curves in Galois $\ell$-extensions

### v1

@@ 2.3. The method of Reichardt and Scholz
Next, we consider the case in which $p$ is ramified. Note that $p\equiv 1\pmod{\ell^{N}}$ by the property $(\mathfrak{S}_{N})$ and thus $p\neq\ell$ and the ramification of $p$ in $L$ is tame. Choose a prime $v$ of $L$ which lies above $p$. We have that $(\mathcal{O}_{L_{v}}/v)^{\times}=(\mathbb{Z}/p\mathbb{Z})^{\times}$. Choose a uniformizer $\pi$ of $L_{v}$ and consider the homomorphism $\lambda:G_{p}\to(\mathbb{Z}/p\mathbb{Z})^{\times}$ defined by $\sigma\mapsto\overline{\left(\frac{\sigma\pi}{\pi}\right)}$. Since $\mathbb{Z}/p\mathbb{Z}$ is the residue field of $L_{v}$, it is easy to check that $\lambda$ is a homomorphism. The kernel of $\lambda$ is the wild inertia subgroup of $G_{p}$, which is trivial. Thus, $G_{p}$ injects into $(\mathbb{Z}/p\mathbb{Z})^{\times}$. It follows therefore that $\varphi_{p}$ factors through a map $\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow\operatorname{G}_{p}$, where $F$ is the maximal abelian tamely ramified extension of $\mathbb{Q}_{p}$ with exponent dividing $\ell^{N}$. More explicitly, $F=F_{1}\cdot F_{2}$, where $F_{1}$ is the unramified extension of $\mathbb{Q}_{p}$ of degree $\ell^{N}$ and $F_{2}:=\mathbb{Q}_{p}(p^{1/\ell^{N}})$. Because $G_{p}$ is abelian, it follows (from the same argument as in the unramified case) that $\widetilde{G}_{p}$ is abelian. Note that $\operatorname{Gal}(F/\mathbb{Q}_{p})\simeq(\mathbb{Z}/\ell^{N}\mathbb{Z})\oplus(\mathbb{Z}/\ell^{N}\mathbb{Z})$. As $G_{p}$ is a quotient of $\operatorname{Gal}(F/\mathbb{Q}_{p})$, we find that $G_{p}\simeq\mathbb{Z}/\ell^{k}\mathbb{Z}$, or $G_{p}\simeq\mathbb{Z}/\ell^{k_{1}}\mathbb{Z}\oplus\mathbb{Z}/\ell^{k_{2}}\mathbb{Z}$ for some $k_{1},k_{2}\geq 1$. Since $\widetilde{G}_{p}$ is abelian and fits into a non-split sequence (2.4), we find that $\widetilde{G}_{p}\simeq\mathbb{Z}/\ell^{k+1}\mathbb{Z}$ or $\widetilde{G}_{p}\simeq\mathbb{Z}/\ell^{k_{1}+1}\mathbb{Z}\oplus\mathbb{Z}/\ell^{k_{2}}\mathbb{Z}$ respectively. In both cases it is clear that $\varphi_{p}:\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow G_{p}$ lifts to a homomorphism $\widetilde{\varphi}_{p}:\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow\widetilde{G}_{p}$. This completes the proof. ∎
@@
There is a surjective lift $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ of $\varphi:\operatorname{G}_{\mathbb{Q}}\rightarrow G$.
@@
Proposition 2.8 implies that for each prime $p$, the local embedding problem is solvable, i.e., there is a lift $\widetilde{\varphi}_{p}$ of $\varphi_{p}$. Thus it follows from Proposition 2.7 that a lift $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ of $\varphi:\operatorname{G}_{\mathbb{Q}}\rightarrow G$ exists. Since the extension
@@
The action of $\operatorname{G}_{\mathbb{Q}}$ on $\mathbb{Z}/\ell\mathbb{Z}$ is trivial, allowing us to identify $H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ with $\operatorname{Hom}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$. Let $\widetilde{L}$ and $\widetilde{L}^{\prime}$ be extensions of $L$ that solve the embedding problem, and let $\widetilde{\varphi},\widetilde{\psi}:\operatorname{G}_{\mathbb{Q}}\to\widetilde{G}$ be the corresponding lifts of $\varphi:\operatorname{G}_{\mathbb{Q}}\to G$ respectively. Then for $\sigma\in\operatorname{G}_{\mathbb{Q}}$, we have that
@@
The map $f:=\widetilde{\psi}\widetilde{\varphi}^{-1}:\operatorname{G}_{\mathbb{Q}}\rightarrow\mathbb{Z}/\ell\mathbb{Z}$ is a homomorphism. In greater detail,
$\begin{split}f(\sigma_{1}\sigma_{2})=&\widetilde{\psi}(\sigma_{1}\sigma_{2})\widetilde{\varphi}(\sigma_{1}\sigma_{2})^{-1}\\ =&\widetilde{\psi}(\sigma_{1})\widetilde{\psi}(\sigma_{2})\widetilde{\varphi}(\sigma_{2})^{-1}\widetilde{\varphi}(\sigma_{1})^{-1}\\ =&\widetilde{\psi}(\sigma_{1})\widetilde{\varphi}(\sigma_{1})^{-1}\widetilde{\psi}(\sigma_{2})\widetilde{\varphi}(\sigma_{2})^{-1}=f(\sigma_{1})f(\sigma_{2})\\ \end{split}$
wherein the second last equality follows since $\widetilde{\psi}(\sigma_{2})\widetilde{\varphi}(\sigma_{2})^{-1}$ belongs to the center of $G$. It is clear that $f(1)=1$ and $f(\sigma^{-1})=f(\sigma)^{-1}$. Conversely, suppose that $f\in H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ and $\widetilde{L}/L/\mathbb{Q}$ solves the embedding problem, with $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ the corresponding homomorphism. Then, $\widetilde{\psi}:=\widetilde{\varphi}f:\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ is a well defined surjective homomorphism that lifts $\varphi$. This also gives rise to a solution $\widetilde{L}^{\prime}/L/K$ of the embedding problem.
@@
$\left\{\widetilde{L}^{f}/L/\mathbb{Q}:\widetilde{L}^{f}\text{ solves the embedding problem}\right\}\leftrightarrow H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z}).$
Let $C$ be a finite dimensional vector space over $\mathbb{Z}/\ell\mathbb{Z}$ equipped with an action of $\operatorname{G}_{\mathbb{Q}}$. Set $C^{*}:=\operatorname{Hom}\left(C,\mu_{\ell}\right)$ with the induced Galois module structure. Given a vector space $V$ over $\mathbb{Z}/\ell\mathbb{Z}$, denote by $V^{\vee}:=\operatorname{Hom}(V,\mathbb{Z}/\ell\mathbb{Z})$. A prime $p$ is said to be ramified in $C$ if the inertia group $\operatorname{I}_{p}\subset\operatorname{G}_{\mathbb{Q}_{p}}$ acts non-trivially on $C$. The field $\mathbb{Q}(C)$ is the Galois extension of $\mathbb{Q}$ cut out by $C$. In greater detail, $\mathbb{Q}(C):=\overline{\mathbb{Q}}^{\operatorname{ker}\rho_{C}}$ where $\rho_{C}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}(C)$ is the representation of $\operatorname{G}_{\mathbb{Q}}$ on $C$. Note that a prime $p$ is ramified in $C$ if and only if it ramifies in $\mathbb{Q}(C)$. Let $\Sigma_{C}$ be the set of primes $p$ such that either $p\neq\ell$ and $p$ is ramified in $C$, or $p=\ell$. Fix a finite set of primes $S$ containing $\Sigma_{C}$ and let $\mathbb{Q}_{S}$ be the maximal algebraic extension of $\mathbb{Q}$ in which all primes $p\notin S$ are unramified. Given a number field $F\subset\mathbb{Q}_{S}$, we set $H^{i}(\mathbb{Q}_{S}/F,\cdot):=H^{i}(\operatorname{Gal}(\mathbb{Q}_{S}/F),\cdot)$ where $i=1,2$. Moreover, we set $H^{i}(\mathbb{Q}_{p},\cdot):=H^{i}(\operatorname{G}_{\mathbb{Q}_{p}},\cdot)$. We define:
(${}^{i}_{S}(C):=\operatorname{ker}\left(H^{i}(\mathbb{Q}_{S}/\mathbb{Q},C)\longrightarrow\bigoplus_{p\in S}H^{i}(\mathbb{Q}_{p},C)\right).$
Global duality for -groups [NSW08, Theorem 8.6.7] states that there is a natural isomorphism:
(${}^{2}_{S}(C)\simeq{}^{1}_{S}(C^{*})^{\vee}.$ (2.5)
@@
$\langle\cdot,\cdot\rangle_{p}:H^{1}(\mathbb{Q}_{p},C)\times H^{1}(\mathbb{Q}_{p},C^{*})\xrightarrow{\cup}H^{1}(\mathbb{Q}_{p},\mu_{\ell})\xrightarrow{\sim}\mathbb{Z}/\ell\mathbb{Z}.$
@@
$\mathcal{L}_{p}^{\perp}:=\{v\in H^{1}(\mathbb{Q}_{p},C^{*})\mid\langle w,v\rangle_{p}=0\text{ for all }w\in\mathcal{L}_{p}\}.$
@@
$\begin{split}&H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},C):=\operatorname{ker}\left\{H^{1}(\mathbb{Q}_{S}/\mathbb{Q},C)\longrightarrow\bigoplus_{p\in S}\frac{H^{1}(\mathbb{Q}_{p},C)}{\mathcal{L}_{p}}\right\},\\ &H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},C^{*}):=\operatorname{ker}\left\{H^{1}(\mathbb{Q}_{S}/\mathbb{Q},C^{*})\longrightarrow\bigoplus_{p\in S}\frac{H^{1}(\mathbb{Q}_{p},C^{*})}{\mathcal{L}_{p}^{\perp}}\right\}.\\ \end{split}$
@@
see for instance, [Tay03, p. 555, l.7]. We set $I_{p}$ to be the inertia subgroup of $\operatorname{G}_{\mathbb{Q}_{p}}$ and given a $\operatorname{G}_{\mathbb{Q}_{p}}$-module $M$, set:
$H^{1}_{\operatorname{nr}}(\mathbb{Q}_{p},M):=\operatorname{image}\{H^{1}(\operatorname{G}_{\mathbb{Q}_{p}}/\operatorname{I}_{p},M^{\operatorname{I}_{p}})\xrightarrow{\operatorname{inf}}H^{1}(\mathbb{Q}_{p},M)\}.$
Consider the Selmer condition $\mathcal{L}^{\operatorname{nr}}$ where $\mathcal{L}_{p}^{\operatorname{nr}}:=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{p},C)$ for all $p\in S$. We note that
@@
where we identify $\operatorname{G}_{\mathbb{Q}_{p}}/\operatorname{I}_{p}$ with $\widehat{\mathbb{Z}}$ (see also [Ram02, Lemma 3]).
@@
Since the Galois action on $C$ is trivial, $H^{1}(\mathbb{Q}_{S}/\mathbb{Q},C)$ consists of homomorphisms $f:\operatorname{G}_{\mathbb{Q}}\rightarrow C$ which are unramified away from $S$, and $H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)$ is the subset of homomorphisms which are unramified at all primes. Such unramified homomorphisms cut out unramified abelian extensions of $\mathbb{Q}$. Since there are no proper abelian unramified extensions of $\mathbb{Q}$, it follows that $H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)=0$. Thus, $\dim H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)=0$. Recall from (2.8) that for $p\in S$,
@@
It follows from Proposition 2.9 that there exists $\widetilde{L}/L/\mathbb{Q}$ solving the embedding problem. Let $\varphi:\operatorname{G}_{\mathbb{Q}}\rightarrow G$ and $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ be the homomorphisms corresponding to $L$ and $\widetilde{L}$ respectively. Let $\Sigma$ (resp. $S$) be the set of primes which ramify in $L$ (resp. $\widetilde{L}$). We modify $\widetilde{\varphi}$ by a class $f\in H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ such that $\widetilde{\psi}:=\widetilde{\varphi}f$ is unramified at all primes $p\notin\Sigma$. For each prime $p\in S$ we can choose a lift $\widetilde{\psi}_{p}:\operatorname{G}_{\mathbb{Q}_{p}}\rightarrow\widetilde{G}_{p}$ of $\varphi_{p}$ such that for $p\in S\setminus\Sigma$, $\widetilde{\psi}_{p}$ is unramified. Such a lift can be constructed as in the proof of Proposition 2.8. For $p\in S\setminus\Sigma$, let $f_{p}\in H^{1}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$ be such that $\widetilde{\psi}_{p}=\widetilde{\varphi}_{p}f_{p}$. On the other hand, for $p\in\Sigma$, set $f_{p}:=0$. Consider the tuple of elements
@@
By part 2 of Lemma 2.12, there exists $f\in H^{1}(\mathbb{Q}_{S}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})$ which restricts to $(f_{p})_{p\in S}$. Note that by construction, $f_{|\operatorname{I}_{p}}=f_{p|\operatorname{I}_{p}}$ for all $p\in S$. It follows that $\widetilde{\psi}:=\widetilde{\varphi}f$ is unramified at all primes $p\in S\setminus\Sigma$. Let $\widetilde{L}^{\prime}$ be the extension corresponding to $\widetilde{\psi}$, then $\widetilde{L}^{\prime}/L/\mathbb{Q}$ is a solution to the embedding problem which is unramified away from $\Sigma$. ∎
@@
[…]
@@
[…]
@@
Now, fix $p=p_{i}$ and let $I_{p}^{\prime}$ (resp. $G_{p}^{\prime}$) is the inertia (resp. decomposition) group of $p$ in $\widetilde{L}$. On the other hand, we set $\widetilde{I}_{p}$ (resp. $\widetilde{G}_{p}$) to be the inverse image of $I_{p}$ (resp. $G_{p}$) with respect to the map $\pi:\widetilde{G}\rightarrow G$. Since $I_{p}=G_{p}$, it follows that $\widetilde{I}_{p}=\widetilde{G}_{p}$. Since the ramification of $p$ is tame, the inertia groups $I_{p}$ and $I_{p}^{\prime}$ are cyclic. Suppose that $I_{p}\simeq\mathbb{Z}/\ell^{\alpha}\mathbb{Z}$. From the proof of Proposition 2.8, $\widetilde{G}_{p}$ is abelian, hence so is $\widetilde{I}_{p}$. Therefore, there are two possibilities for $\widetilde{I}_{p}$.
@@
Since the map $I_{p}^{\prime}\rightarrow I_{p}$ is surjective, it follows that $I_{p}^{\prime}=\widetilde{I}_{p}=\widetilde{G}_{p}$. This forces $I_{p}^{\prime}=G_{p}^{\prime}$. In this case, we simply set $c_{p}:=1$.

### v2

@@ 2.3. The method of Reichardt and Scholz
Next, we consider the case in which $p$ is ramified. Note that $p\equiv 1\pmod{\ell^{N}}$ by the property $(\mathfrak{S}_{N})$ and thus $p\neq\ell$ and the ramification of $p$ in $L$ is tame. Choose a prime $v$ of $L$ which lies above $p$. We have that $(\mathcal{O}_{L_{v}}/v)^{\times}=(\mathbb{Z}/p\mathbb{Z})^{\times}$. Choose a uniformizer $\pi$ of $L_{v}$ and consider the homomorphism $\lambda\mathrel{\mathop{\ordinarycolon}}G_{p}\to(\mathbb{Z}/p\mathbb{Z})^{\times}$ defined by $\sigma\mapsto\overline{\left(\frac{\sigma\pi}{\pi}\right)}$. Since $\mathbb{Z}/p\mathbb{Z}$ is the residue field of $L_{v}$, it is easy to check that $\lambda$ is a homomorphism. The kernel of $\lambda$ is the wild inertia subgroup of $G_{p}$, which is trivial. Thus, $G_{p}$ injects into $(\mathbb{Z}/p\mathbb{Z})^{\times}$. It follows therefore that $\varphi_{p}$ factors through a map $\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow\operatorname{G}_{p}$, where $F$ is the maximal abelian tamely ramified extension of $\mathbb{Q}_{p}$ with exponent dividing $\ell^{N}$. More explicitly, $F=F_{1}\cdot F_{2}$, where $F_{1}$ is the unramified extension of $\mathbb{Q}_{p}$ of degree $\ell^{N}$ and $F_{2}\mathrel{\mathop{\ordinarycolon}}=\mathbb{Q}_{p}(p^{1/\ell^{N}})$. Because $G_{p}$ is abelian, it follows (from the same argument as in the unramified case) that $\widetilde{G}_{p}$ is abelian. Note that $\operatorname{Gal}(F/\mathbb{Q}_{p})\simeq(\mathbb{Z}/\ell^{N}\mathbb{Z})\oplus(\mathbb{Z}/\ell^{N}\mathbb{Z})$. As $G_{p}$ is a quotient of $\operatorname{Gal}(F/\mathbb{Q}_{p})$, we find that $G_{p}\simeq\mathbb{Z}/\ell^{k}\mathbb{Z}$, or $G_{p}\simeq\mathbb{Z}/\ell^{k_{1}}\mathbb{Z}\oplus\mathbb{Z}/\ell^{k_{2}}\mathbb{Z}$ for some $k_{1},k_{2}\geq 1$. Since $\widetilde{G}_{p}$ is abelian and fits into a non-split sequence (2.4), we find that $\widetilde{G}_{p}\simeq\mathbb{Z}/\ell^{k+1}\mathbb{Z}$ or $\widetilde{G}_{p}\simeq\mathbb{Z}/\ell^{k_{1}+1}\mathbb{Z}\oplus\mathbb{Z}/\ell^{k_{2}}\mathbb{Z}$ respectively. In both cases it is clear that $\varphi_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow G_{p}$ lifts to a homomorphism $\widetilde{\varphi}_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(F/\mathbb{Q}_{p})\rightarrow\widetilde{G}_{p}$. This completes the proof. ∎
@@
There is a surjective lift $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\twoheadrightarrow G$ of $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow G$.
@@
Proposition 2.8 implies that for each prime $p$, the local embedding problem is solvable, i.e., there is a lift $\widetilde{\varphi}_{p}$ of $\varphi_{p}$. Thus it follows from Proposition 2.7 that a lift $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ of $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow G$ exists. Since the extension
@@
The action of $\operatorname{G}_{\mathbb{Q}}$ on $\mathbb{Z}/\ell\mathbb{Z}$ is trivial, allowing us to identify $H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ with $\operatorname{Hom}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$. Let $\widetilde{L}$ and $\widetilde{L}^{\prime}$ be extensions of $L$ that solve the embedding problem, and let $\widetilde{\varphi},\widetilde{\psi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to\widetilde{G}$ be the corresponding lifts of $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to G$ respectively. Then for $\sigma\in\operatorname{G}_{\mathbb{Q}}$, we have that
@@
The map $f\mathrel{\mathop{\ordinarycolon}}=\widetilde{\psi}\widetilde{\varphi}^{-1}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\mathbb{Z}/\ell\mathbb{Z}$ is a homomorphism. In greater detail,
[…]
wherein the second last equality follows since $\widetilde{\psi}(\sigma_{2})\widetilde{\varphi}(\sigma_{2})^{-1}$ belongs to the center of $G$. It is clear that $f(1)=1$ and $f(\sigma^{-1})=f(\sigma)^{-1}$. Conversely, suppose that $f\in H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ and $\widetilde{L}/L/\mathbb{Q}$ solves the embedding problem, with $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ the corresponding homomorphism. Then, $\widetilde{\psi}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\varphi}f\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ is a well defined surjective homomorphism that lifts $\varphi$. This also gives rise to a solution $\widetilde{L}^{\prime}/L/K$ of the embedding problem.
@@
$\left\{\widetilde{L}^{f}/L/\mathbb{Q}\mathrel{\mathop{\ordinarycolon}}\widetilde{L}^{f}\text{ solves the embedding problem}\right\}\leftrightarrow H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z}).$
Let $C$ be a finite dimensional vector space over $\mathbb{Z}/\ell\mathbb{Z}$ equipped with an action of $\operatorname{G}_{\mathbb{Q}}$. Set $C^{*}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Hom}\left(C,\mu_{\ell}\right)$ with the induced Galois module structure. Given a vector space $V$ over $\mathbb{Z}/\ell\mathbb{Z}$, denote by $V^{\vee}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Hom}(V,\mathbb{Z}/\ell\mathbb{Z})$. A prime $p$ is said to be ramified in $C$ if the inertia group $\operatorname{I}_{p}\subset\operatorname{G}_{\mathbb{Q}_{p}}$ acts non-trivially on $C$. The field $\mathbb{Q}(C)$ is the Galois extension of $\mathbb{Q}$ cut out by $C$. In greater detail, $\mathbb{Q}(C)\mathrel{\mathop{\ordinarycolon}}=\overline{\mathbb{Q}}^{\operatorname{ker}\rho_{C}}$ where $\rho_{C}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}(C)$ is the representation of $\operatorname{G}_{\mathbb{Q}}$ on $C$. Note that a prime $p$ is ramified in $C$ if and only if it ramifies in $\mathbb{Q}(C)$. Let $\Sigma_{C}$ be the set of primes $p$ such that either $p\neq\ell$ and $p$ is ramified in $C$, or $p=\ell$. Fix a finite set of primes $S$ containing $\Sigma_{C}$ and let $\mathbb{Q}_{S}$ be the maximal algebraic extension of $\mathbb{Q}$ in which all primes $p\notin S$ are unramified. Given a number field $F\subset\mathbb{Q}_{S}$, we set $H^{i}(\mathbb{Q}_{S}/F,\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{Gal}(\mathbb{Q}_{S}/F),\cdot)$ where $i=1,2$. Moreover, we set $H^{i}(\mathbb{Q}_{p},\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{\mathbb{Q}_{p}},\cdot)$. We define:
$^{i}_{S}(C)\mathrel{\mathop{\ordinarycolon}}=\operatorname{ker}\left(H^{i}(\mathbb{Q}_{S}/\mathbb{Q},C)\longrightarrow\bigoplus_{p\in S}H^{i}(\mathbb{Q}_{p},C)\right).$
Global duality for -groups [NSW, Theorem 8.6.7] states that there is a natural isomorphism:
[…]
@@
$\langle\cdot,\cdot\rangle_{p}\mathrel{\mathop{\ordinarycolon}}H^{1}(\mathbb{Q}_{p},C)\times H^{1}(\mathbb{Q}_{p},C^{*})\xrightarrow{\cup}H^{1}(\mathbb{Q}_{p},\mu_{\ell})\xrightarrow{\sim}\mathbb{Z}/\ell\mathbb{Z}.$
@@
$\mathcal{L}_{p}^{\perp}\mathrel{\mathop{\ordinarycolon}}=\{v\in H^{1}(\mathbb{Q}_{p},C^{*})\mid\langle w,v\rangle_{p}=0\text{ for all }w\in\mathcal{L}_{p}\}.$
@@
[…]
@@
see for instance, [tayloricosahedral, p. 555, l.7]. We set $I_{p}$ to be the inertia subgroup of $\operatorname{G}_{\mathbb{Q}_{p}}$ and given a $\operatorname{G}_{\mathbb{Q}_{p}}$-module $M$, set:
$H^{1}_{\operatorname{nr}}(\mathbb{Q}_{p},M)\mathrel{\mathop{\ordinarycolon}}=\operatorname{image}\{H^{1}(\operatorname{G}_{\mathbb{Q}_{p}}/\operatorname{I}_{p},M^{\operatorname{I}_{p}})\xrightarrow{\operatorname{inf}}H^{1}(\mathbb{Q}_{p},M)\}.$
Consider the Selmer condition $\mathcal{L}^{\operatorname{nr}}$ where $\mathcal{L}_{p}^{\operatorname{nr}}\mathrel{\mathop{\ordinarycolon}}=H^{1}_{\operatorname{nr}}(\mathbb{Q}_{p},C)$ for all $p\in S$. We note that
@@
where we identify $\operatorname{G}_{\mathbb{Q}_{p}}/\operatorname{I}_{p}$ with $\widehat{\mathbb{Z}}$ (see also [raviFM, Lemma 3]).
@@
Since the Galois action on $C$ is trivial, $H^{1}(\mathbb{Q}_{S}/\mathbb{Q},C)$ consists of homomorphisms $f\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow C$ which are unramified away from $S$, and $H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)$ is the subset of homomorphisms which are unramified at all primes. Such unramified homomorphisms cut out unramified abelian extensions of $\mathbb{Q}$. Since there are no proper abelian unramified extensions of $\mathbb{Q}$, it follows that $H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)=0$. Thus, $\dim H^{1}_{\mathcal{L}^{\operatorname{nr}}}(\mathbb{Q}_{S}/\mathbb{Q},C)=0$. Recall from (2.8) that for $p\in S$,
@@
It follows from Proposition 2.9 that there exists $\widetilde{L}/L/\mathbb{Q}$ solving the embedding problem. Let $\varphi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow G$ and $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\widetilde{G}$ be the homomorphisms corresponding to $L$ and $\widetilde{L}$ respectively. Let $\Sigma$ (resp. $S$) be the set of primes which ramify in $L$ (resp. $\widetilde{L}$). We modify $\widetilde{\varphi}$ by a class $f\in H^{1}(\operatorname{G}_{\mathbb{Q}},\mathbb{Z}/\ell\mathbb{Z})$ such that $\widetilde{\psi}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\varphi}f$ is unramified at all primes $p\notin\Sigma$. For each prime $p\in S$ we can choose a lift $\widetilde{\psi}_{p}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}_{p}}\rightarrow\widetilde{G}_{p}$ of $\varphi_{p}$ such that for $p\in S\setminus\Sigma$, $\widetilde{\psi}_{p}$ is unramified. Such a lift can be constructed as in the proof of Proposition 2.8. For $p\in S\setminus\Sigma$, let $f_{p}\in H^{1}(\mathbb{Q}_{p},\mathbb{Z}/\ell\mathbb{Z})$ be such that $\widetilde{\psi}_{p}=\widetilde{\varphi}_{p}f_{p}$. On the other hand, for $p\in\Sigma$, set $f_{p}\mathrel{\mathop{\ordinarycolon}}=0$. Consider the tuple of elements
@@
By part 2 of Lemma 2.12, there exists $f\in H^{1}(\mathbb{Q}_{S}/\mathbb{Q},\mathbb{Z}/\ell\mathbb{Z})$ which restricts to $(f_{p})_{p\in S}$. Note that by construction, $f_{|\operatorname{I}_{p}}=f_{p|\operatorname{I}_{p}}$ for all $p\in S$. It follows that $\widetilde{\psi}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\varphi}f$ is unramified at all primes $p\in S\setminus\Sigma$. Let $\widetilde{L}^{\prime}$ be the extension corresponding to $\widetilde{\psi}$, then $\widetilde{L}^{\prime}/L/\mathbb{Q}$ is a solution to the embedding problem which is unramified away from $\Sigma$. ∎
@@
    FLFL and $\mathbb{Q}(\mu_{\ell},\sqrt[\ell]{p_{1}},\dots,\sqrt[\ell]{p_{m}})$ are linearly disjoint.
@@
The prime $\ell$ is unramified in $L$ and is totally ramified in $F$. Therefore, $F$ and $L$ are linearly disjoint. The Galois group $\operatorname{Gal}(\mathbb{Q}(\mu_{\ell},\sqrt[\ell]{p_{1}},\dots,\sqrt[\ell]{p_{m}})/\mathbb{Q})$ is isomorphic to $(\mathbb{Z}/\ell\mathbb{Z})^{m}\rtimes(\mathbb{Z}/\ell\mathbb{Z})^{\times}$, where multiplication is given by $(x,y)\cdot(x^{\prime},y^{\prime})=(x+yx^{\prime},yy^{\prime})$. It is easy to see that this group has no quotient of order $\ell$. On the other hand, $FL$ is an $\ell$-group. Hence we find that $FL\,\cap\,\mathbb{Q}(\mu_{\ell},\sqrt[\ell]{p_{1}},\dots,\sqrt[\ell]{p_{m}})=\mathbb{Q}$. ∎
@@
Now, fix $p=p_{i}$ and let $I_{p}^{\prime}$ (resp. $G_{p}^{\prime}$) is the inertia (resp. decomposition) group of $p$ in $\widetilde{L}$. On the other hand, we set $\widetilde{I}_{p}$ (resp. $\widetilde{G}_{p}$) to be the inverse image of $I_{p}$ (resp. $G_{p}$) with respect to the map $\pi\mathrel{\mathop{\ordinarycolon}}\widetilde{G}\rightarrow G$. Since $I_{p}=G_{p}$, it follows that $\widetilde{I}_{p}=\widetilde{G}_{p}$. Since the ramification of $p$ is tame, the inertia groups $I_{p}$ and $I_{p}^{\prime}$ are cyclic. Suppose that $I_{p}\simeq\mathbb{Z}/\ell^{\alpha}\mathbb{Z}$. From the proof of Proposition 2.8, $\widetilde{G}_{p}$ is abelian, hence so is $\widetilde{I}_{p}$. Therefore, there are two possibilities for $\widetilde{I}_{p}$.
@@
Since the map $I_{p}^{\prime}\rightarrow I_{p}$ is surjective, it follows that $I_{p}^{\prime}=\widetilde{I}_{p}=\widetilde{G}_{p}$. This forces $I_{p}^{\prime}=G_{p}^{\prime}$. In this case, we simply set $c_{p}\mathrel{\mathop{\ordinarycolon}}=1$.